the angle of a triangle are x°, (x-10°) and 90°. find the angle of the triangle topic:basic geometric theorems
Answer by Boreal(15235)

Angles of a triangle add to 180 degrees.
So
180=x + (x-10) +90 = 2x -10 +90 = 2x +80
subtract 80 from each side
100= 2x
x=50 degrees
x-10=40 degrees
50 +40= 90
The right angle is 90
90 deg + 90 deg =180 deg
check